OpenStudy (anonymous):

How would you figure out if a binomial is a factor of a polynomial function?

OpenStudy (lgbasallote):

divide the polynomial function by the binomial. If the remainder is 0, then the binomial is a factor of the polynomial function

OpenStudy (anonymous):

Oh! Okay!!

OpenStudy (lgbasallote):

or you can use the remainder theorem given \[f(x) \div (x-a)\] x- a is a factor IF \[f(a) = 0\]

OpenStudy (anonymous):

So my question was... Is the binomial (x + 3) a factor of: x^3+4x^2+x-6? i would say yes because i divided i together and the remainder is 0 is this correct?

OpenStudy (lgbasallote):

= (-3)^3 + 4(-3)^2 + (-3) - 6 = -27 + 4(9) - 3 - 6 = -27 + 36 - 3 - 6 = -36 + 36 = 0 yup..it's a factor

OpenStudy (anonymous):

awesome! I feel really proud of myself :D Thanks again!

OpenStudy (lgbasallote):

welcome